need formula
Your bank account was overdrawn by $14. You then were charged a $35 overdraft fee. The bank applies a $5 fee for each day your account is over drawn. Your account is overdrawn for 3 days when you deposit a check for $256. What is your account balance after depositing the check? Will you have enough money to pay your $52 electric bill (assuming no other money is spent)?

Found 2 solutions by ReadingBoosters, jim_thompson5910:
Answer by ReadingBoosters(3246)

x = balance after depositing check
-14 + -35 + -5(3) + 256 = x
If x is greater than 52, then yes there will be enough to cover electric bill
If x is less than 52, then no there will not be enought to cover electric bill

Answer by jim_thompson5910(35256)

"Your bank account was overdrawn by $14"

means that you're $14 in the hole. So the balance is at -14 dollars.

"You then were charged a $35 overdraft fee"

means you're an additional $35 in the hole. So the balance is now at -14-35 = -49 dollars.

Let x = number of days the account is over drawn

If a $5 fee is applied each day the account is over drawn, then you're subtracting 5 dollars each day. So over the course of x days, you will have subtracted 5x dollars total.

So the final balance after x days is -49 - 5x

After 3 days, the final balance will be...

-49 - 5x

-49 - 5(3)

-49 - 15

-64

So after 3 days, the final balance will be -64 dollars.

Now add on the $256 from the check you deposit: -64+256 = 192

Since $192 is larger than $52, this means you will have enough money to pay your $52 electric bill (assuming no other money is spent).